EMISSION SPECTROSCOPY FOLLOWING THE MULTIPHOTON PHOTOLYSIS OF HALOMETHANES AT NEAR-ULTRAVIOLET WAVELENGTHS

CHEN-NAN LIU, HSIANG-FU LIAO, GUANG-YI HOU, SHI-XING YANG, and \underline{BOR-CHEN CHANG}, Department of Chemistry, National Central University, Jhongli 32001, Taiwan.

Emission spectroscopy including nascent emission and laser-induced dispersed fluorescence was adopted to investigate the multiphoton photolysis mechanism of halomethanes at near-ultraviolet wavelengths in a slow flow system. In the 266 nm photolysis of the interested halomethanes (CHBr_{3}, CHBr_{2}Cl, CHBrCl_{2}, CH_{2}Br_{2}, CHI_{3}, CH_{2}I_{2}, and CH_{3}I), several excited species such as CH (A^{2}Δ, B^{2}Σ^{−}, and C^{2}Σ^{+}), atomic Br or I, and C_{2} (d^{3}Π_{g}) were observed in the nascent emission spectra. Halomethylenes (CHX, X= Br, Cl, I), the reactive intermediates, were not observed in nascent emission spectra, but they can be found using laser-induced dispersed fluorescence spectroscopy following excitation of their Ã^{1}A′′(0,v_{2},0)←X̃^{1}A′(0,0,0) transitions. Interestingly, CHBr was seen only in the photolysis of CHBr_{3}, whereas CHCl was only discovered when the precursor is CHBr_{2}Cl or CHBrCl_{2}. The photolysis laser power dependence and emission waveform measurements were also conducted. In addition, nascent emission spectra following the photolysis at longer near-ultraviolet wavelengths (280 nm and 355 nm) were acquired. The results show the distinctive differences between the photolysis of bromomethanes (CHBr_{3}, CHBr_{2}Cl, CHBrCl_{2}, and CH_{2}Br_{2}) and that of iodomethanes (CHI_{3}, CH_{2}I_{2}, and CH_{3}I). Our recent progress will be presented.